# COOD: Concept-based Zero-shot OOD Detection

### v1

@@ Abstract
How can models effectively detect out-of-distribution (OOD) samples in complex, multi-label settings without extensive retraining? Existing OOD detection methods struggle to capture the intricate semantic relationships and label co-occurrences inherent in multi-label settings, often requiring large amounts of training data and failing to generalize to unseen label combinations. While large language models have revolutionized zero-shot OOD detection, they primarily focus on single-label scenarios, leaving a critical gap in handling real-world tasks where samples can be associated with multiple interdependent labels. To address these challenges, we introduce COOD, a novel zero-shot multi-label OOD detection framework. COOD leverages pre-trained vision-language models, enhancing them with a concept-based label expansion strategy and a new scoring function. By enriching the semantic space with both positive and negative concepts for each label, our approach models complex label dependencies, precisely differentiating OOD samples without the need for additional training. Extensive experiments demonstrate that our method significantly outperforms existing approaches, achieving approximately 95% average AUROC on both VOC and COCO datasets, while maintaining robust performance across varying numbers of labels and different types of OOD samples.
1 Introduction
@@
As machine learning (ML) models become essential in a range of real-world applications, out-of-distribution (OOD) detection has gained increasing importance [36]. OOD detection is particularly critical in fields such as autonomous driving [3], medical diagnostics [10], and surveillance [29], where managing data that deviates from the training distribution—known as in-distribution (ID)—is necessary to avoid safety risks or incorrect decisions [9]. Thus, developing robust OOD detection techniques is key to ensuring model reliability in unpredictable environments.
@@
Large Models for OOD Detection (LM-OOD). The rise of large models, particularly Vision-Language Models (VLMs) and MultiModal Large Language Models (MLLMs), has redefined the landscape of OOD detection [34]. Traditional OOD methods, such as Maximum Softmax Probability (MSP) and Mahalanobis distance-based techniques [15], typically require extensive task-specific training on in-distribution (ID) data. In contrast, LLM-based methods leverage pre-trained knowledge, allowing for zero-shot and few-shot OOD detection capabilities. Models like CLIP [27] exemplify this shift by achieving effective OOD detection with minimal task-specific training, relying on robust, pre-trained representations from multimodal datasets. Recent advancements, such as NegLabel [11], refine zero-shot OOD detection by introducing negative mining strategies to identify semantically meaningful OOD categories. This approach highlights that carefully selected negative labels can significantly improve detection performance without retraining. LM-OOD methods offer distinct advantages over traditional approaches, including strong performance in limited data, high adaptability across various tasks, and improved computational efficiency through reduced dependence on task-specific data preparation and retraining [22, 24, 34].
@@
Figure 1: Motivation for Concept-Based OOD Detection. Traditional methods struggle with complex multi-label cases. Our approach expands the label space with positive and negative concepts, enabling robust detection of complex OOD samples like “Okapi" and “Spork."
@@
Limitations in Current LM-OOD. Large models for OOD detection are primarily designed for single-label tasks, where each input corresponds to a single, definitive label [6, 22]. However, many real-world applications are inherently multi-label in nature: medical imaging often requires identifying multiple coexisting conditions [12, 10]. Current state-of-the-art methods like NegLabel [11] encounter limitations in these multi-label contexts, as seen in Fig. 1: First, the complexity of semantic similarity computations increases significantly with multiple concurrent labels compared to single labels. Second, these methods assume OOD samples are semantically distant from ID classes, which may not hold in multi-label settings where novel combinations of known concepts could form valid OOD cases [33, 25]. Additionally, these approaches struggle to effectively model complex label co-occurrence patterns and conditional dependencies, both crucial in multi-label contexts [39, 37]. These limitations highlight an essential gap in adapting LM-OOD methods for multi-label tasks, where it is necessary to differentiate between known and unknown label combinations within complex label spaces for practical application [24, 36].
@@
Our Proposal: Extending LM-OOD to Multi-Label Settings. In this work, we introduce COOD (concept-based OOD detection), the first multi-label OOD detection framework that leverages VLM models, specifically CLIP [33], in a zero-shot setting. COOD addresses the challenges of multi-label OOD detection by incorporating a concept-based label expansion strategy. It enriches the base label set with two fine-grained concepts: positive concepts, which capture complex semantic details closely related to ID classes, and negative concepts, which, filtered by a similarity threshold, introduce semantically distant features to strengthen ID-OOD separation. COOD embeds these expanded concept labels alongside the original base labels, creating a refined semantic space that enables precise detection of subtle differences between ID and OOD samples, without additional training. By leveraging CLIP’s extensive pre-trained multimodal knowledge, COOD achieves robust and flexible OOD detection in complex, multi-label scenarios.
Our technical contributions are summarized as follows:
@@
  • Effectiveness: Our method demonstrates superior performance across standard benchmarks, outperforming existing methods in multi-label OOD detection while maintaining efficiency and robustness in real-world settings. Our method achieves an OOD sample detection efficiency of approximately 800 images per second on the CLIP-B/16.
2 Proposed COOD Method
Figure 2: Overview of COOD (§2.2 ). The Concept Generation module (§2.3) uses LLMs to expand base labels into positive ($\mathcal{P}$) and negative ($\mathcal{N}$) concept sets, enhancing the ID-OOD boundary. Positive concepts capture fine-grained, ID-aligned features, while negative concepts provide contrasting OOD-aligned features. The Similarity and ID Score Computation module (§2.4 ) encodes an input image and computes similarity scores with $\mathcal{P}$, $\mathcal{B}$ (base), and $\mathcal{N}$ sets. An ID score based on top-$k$ similarities then classifies the image for ID/OOD.
2.1 Preliminaries on OOD Detection
In multi-label OOD detection, we aim to determine whether a given input image $I\in\mathcal{I}$ belongs to the ID or represents an unknown OOD sample. We denote the set of base labels as =$\mathcal{B}=\{\ell_{1},\ell_{2},\dots,\ell_{|\mathcal{B}|}\}$, representing known classes within the ID. Each image $I$ can be associated with multiple labels from $\mathcal{B}$, capturing the multi-label nature of this task.
To formalize this, we define a scoring function, $S_{\text{ID}}:\mathcal{I}\rightarrow\mathbb{R}$, which assigns an ID score to each image. For simplicity, we frame our approach around this ID score rather than the OOD score, with the following criteria:
@@ 2.1 Preliminaries on OOD Detection
where $\gamma$ is a threshold determined through validation, and $\mathcal{D}_{\text{ID}}$ and $\mathcal{D}_{\text{OOD}}$ are ID and OOD distributions, respectively. In this setup, $S_{\text{ID}}$ provides a measure of similarity to the ID distribution, aiding in distinguishing between known and unknown samples in a multi-label context.
2.1.1 Revisit VLM-based OOD Detection
Vision-Language Models (VLMs) bring powerful multimodal capabilities for OOD detection, allowing more flexible, context-aware identification of OOD samples [24, 34]. Large-scale pre-trained VLMs, such as CLIP [27] and GPT-4V [38], enable models to process images and text prompts together, enhancing adaptability and precision across diverse visual and textual domains [6].
VLM-based OOD detection generally follows two primary approaches [34]: Prompting-based Detection, which directly prompts VLMs to respond with OOD indicators [1], and Contrasting-based Detection, which uses multimodal VLMs pre-trained with contrastive objectives to distinguish OOD samples [22]. This work focuses on the contrasting-based approach in multimodal contexts, as it is well-suited for enhancing OOD detection by amplifying distinctions between ID and OOD classes. Methods like NegLabel [11] and NegPrompt [16] fall under this category. NegLabel enhances OOD detection by introducing negative labels to contrast with in-distribution (ID) classes, while NegPrompt uses learned negative prompts to emphasize OOD differences by contrasting them with ID prompts.
@@ 2.1.1 Revisit VLM-based OOD Detection
where $h$ is the embedding of $x$; $e_{i}$ and $e_{j}^{-}$ are embeddings for ID and negative representations (NegPrompt [16] or NegLabel[11]) from sets $Y$ and $Y^{-}$, respectively; and $\text{sim}(h,e)$ denotes the similarity (e.g., cosine similarity) between embeddings. The numerator measures similarity to ID representations, while the denominator incorporates both ID and negative similarities, amplifying the detection signal. Note that, our definition is a little different, a smaller $S(x)$ means a larger probability of $x$ being an OOD sample.
@@
Moreover, extending single-label OOD detection methods [30, 20] to multi-label contexts presents additional challenges. Multi-label settings often involve significant semantic overlap between labels and co-occurring ID and OOD labels, complicating the extraction of effective negative labels or prompts. These issues highlight the need for advanced detection techniques to capture the complex dependencies and relationships inherent in multi-label OOD detection tasks.
2.2 Overview of COOD Method
The COOD framework improves OOD detection by refining the decision boundary between ID and OOD samples using fine-grained concepts. By introducing positive and negative concept sets, $\mathcal{P}$ and $\mathcal{N}$, COOD leverages LLMs to add contextual information around ID samples, thereby enhancing sensitivity to OOD cases.
COOD is built on two core components (see Fig. 2):
@@ 2.2 Overview of COOD Method
    Concept Generation and Similarity Measure2.3): To distinguish ID from OOD samples, we generate positive concepts ($\mathcal{P}$) and negative concepts ($\mathcal{N}$) that are closely and distantly related to ID classes, respectively. Positive concepts, mined using LLMs, capture domain-relevant features of ID samples at a fine-grained level, while negative concepts are chosen based on their semantic distance from ID classes to enhance contrast.
@@
    Similarity and ID Score Computation2.4): For each input image $I$, we compute similarity scores with these concept sets and the base label set $\mathcal{B}$, forming a robust semantic space for evaluating ID-OOD relationships. After this, we define an ID score $S_{\text{ID}}(I)$ that aggregates the image’s alignment with ID concepts ($\mathcal{B}$ and $\mathcal{P}$) and contrasts it with the negative concepts ($\mathcal{N}$). When this score falls below a predefined threshold $\gamma$, the image is classified as OOD. This scoring method sharpens the decision boundary by leveraging contrasts between positive and negative concepts, giving more accurate OOD detection.
The pseudocode in Algo. 1 outlines the full procedure for COOD. To formalize the OOD detection decision, we compute an ID score and classify an image $X$ as OOD if this score falls below a threshold $\gamma$. Specifically, the decision function $\tilde{Y}$ is defined as follows:
@@
Theoretical Justification. we derive theoretical bounds (§2.5) that validate the scoring function’s effectiveness in assigning higher scores to ID samples than OOD samples, underscoring the robustness of our approach.
Advantages. The proposed COOD approach is computationally efficient, leveraging pre-trained embeddings without requiring additional model training. By incorporating positive and negative concepts, the model gains enhanced semantic understanding, allowing for clearer differentiation between ID and OOD samples. The focus on top-$k$ similarity values makes the method robust to noise, ensuring stable performance across varied datasets and label sets.
Algorithm 1 Multi-Label OOD Detection with COOD
0:  Image $I$, base label set $\mathcal{B}$, threshold $\gamma$, top-$k$ parameter $k$
0:  Classification of $I$ as ID or OOD
1:  Concept Generation2.3): Generate positive concepts $\mathcal{P}$ and negative concepts $\mathcal{N}$ using an LLM.
@@
3:  Top-$k$ Similarity Calculation2.4): Calculate the top-$k$ mean similarity scores, $\mu_{k}(\mathcal{B},I)$, $\mu_{k}(\mathcal{P},I)$, and $\mu_{k}(\mathcal{N},I)$.
4:  ID Score Computation2.4): Using the top-$k$ similarities, compute the ID score $S_{\text{ID}}(I)$ according to Eq. (7).
5:  Decision: if $S_{\text{ID}}(I)>\gamma$ then classify $I$ as ID; else classify $I$ as OOD.
2.3 Concept Generation
Motivation. The concept generation process in COOD creates a rich semantic space to strengthen ID-OOD distinctions. By constructing a set of positive concepts, $\mathcal{P}$, closely aligned with ID samples, and a set of negative concepts, $\mathcal{N}$, that enhances the contrast with OOD samples, we can refine the decision boundary between ID and OOD.
Positive Concept Mining $\mathcal{P}$. To build a comprehensive set of positive concepts that accurately represents ID characteristics, we use a two-stage approach inspired by LF-CBM [26]. This process enriches each base label by capturing specific attributes that reinforce its identity within the ID class.
@@ 2.3 Concept Generation
    Concept Querying: We prompt GPT-4 to generate concepts in three distinct contexts for each target object, using structured prompts tailored to elicit three categories: features (promptF), superclasses (promptS), and commonly associated items (promptC). This contextual querying ensures that each concept captures detailed, complex information, improving the clarity and consistency of the concept pool. Each prompt is crafted to encourage concise, relevant responses without qualifiers, which enhances precision, detailed in the Appendix LABEL:appx:concept_generation.
@@
    Concept Filtering: After generating these candidate concepts, we apply filtering criteria to retain only the most distinctive and relevant features. This step is essential to ensure that the concept set effectively captures the ID characteristics required to differentiate ID from OOD. Our filtering process, detailed in the Appendix LABEL:appx:concept_generation, refines the generated concepts into a cohesive and meaningful set of positive labels.
@@
Negative Concept Mining $\mathcal{N}$. To reinforce the distinction between ID and OOD samples, we develop a set of negative concepts, $\mathcal{N}$, using a process called NegMining [11]. This approach helps us identify concepts that are semantically distant from ID labels, forming a contrasting boundary that enhances OOD detection.
@@
$\mathcal{N}$ =$=\left\{n\mid\text{Sim}(\tilde{n}_{i},\mathcal{B})<\tau_{i}\right\},$ (6)
@@
By expanding the base label set $\mathcal{B}$ to include positive concepts $\mathcal{P}$ and contrasting them with negative concepts $\mathcal{N}$, we create an enriched semantic space. This structure allows for more precise discrimination between ID and OOD samples, capturing both the core characteristics of the ID classes and their semantic boundaries effectively.
2.4 Similarity and ID Score Computation
After generating our positive and negative concepts, COOD computes an ID score for each input image $I$, based on its similarity with the positive concept set $\mathcal{P}$, negative concept set $\mathcal{N}$, and base labels $\mathcal{B}$. This score helps determine if the image aligns more closely with ID or OOD characteristics. ID Score Definition. The ID score $S_{\text{ID}}(I)$ for an image $I$ is:
$S_{\text{ID}}(I)=w_{\mathcal{B}}\cdot\mu_{k}(\mathcal{B},I)-w_{\mathcal{P}}\cdot\mu_{k}(\mathcal{P},I)-w_{\mathcal{N}}\cdot\mu_{k}(\mathcal{N},I)$ (7)
where $w_{\mathcal{B}},w_{\mathcal{P}},$ and $w_{\mathcal{N}}$ are weighting coefficients that determine the relative contribution of each set in computing the ID score, reflecting the importance of the base labels $\mathcal{B}$, positive concepts $\mathcal{P}$, and negative concepts $\mathcal{N}$.
@@ 2.4 Similarity and ID Score Computation
Interpretation of the ID Score. The scoring function $S_{\text{ID}}(I)$ captures the degree of alignment of an image with ID or OOD characteristics. For ID samples, we anticipate high alignment with both the base labels $\mathcal{B}$ and positive concepts $\mathcal{P}$, yielding a high $S_{\text{ID}}(I)$. Conversely, OOD samples will likely align more closely with negative concepts $\mathcal{N}$, resulting in a lower ID score. By focusing on the top-$k$ similarities, this approach enhances robustness to noise and emphasizes the most semantically relevant features.
This mechanism refines the boundary between ID and OOD samples by leveraging positive and negative contrasts, resulting in more accurate and reliable OOD detection.
2.5 Theoretical Justification
Lemma 1 (Separability of ID-OOD).
The scoring function can distinguish between in-distribution (ID) and out-of-distribution (OOD) samples based on their semantic relationships by appropriately selecting weights $w_{P}$ and $w_{\mathcal{N}}$.
Proof.
We assume the following bounds on the similarity scores $\mu_{k}(\cdot,\cdot)$ between an image $I$ and class representations in different contexts when $I$ is either ID or OOD:
$0\leq\mu_{k}(\mathcal{B},I_{\text{OOD}})\leq a<\mu_{k}(\mathcal{B},I_{\text{ID}})\leq A\leq 1,$
$0\leq\mu_{k}(\mathcal{P},I_{\text{OOD}})\leq b<\mu_{k}(\mathcal{P},I_{\text{ID}})\leq B\leq 1,$
$0\leq\mu_{k}(\mathcal{N},I_{\text{ID}})\leq c<\mu_{k}(\mathcal{N},I_{\text{OOD}})\leq C\leq 1$
To establish separability between ID and OOD samples, we examine bounds on the scoring function for in-distribution and out-of-distribution images.
1. Lower Bound for $S_{\text{ID}}$: Based on the assumed bounds, the lower bound of the scoring function $S_{\text{ID}}$ for ID samples is: $S_{\text{ID}}^{\text{lower}}=w_{B}a-w_{P}B-w_{\mathcal{N}}c$.
2. Upper Bound for $S_{\text{ID}}$: The upper bound of the scoring function $S_{\text{ID}}$ for OOD samples is: $S_{\text{ID}}^{\text{upper}}=w_{B}a-w_{\mathcal{N}}C$.
3. Separability Condition: For the scoring function to effectively distinguish between ID and OOD samples, we require: $S_{\text{ID}}^{\text{lower}}-S_{\text{ID}}^{\text{upper}}>0$. Substituting the expressions for $S_{\text{ID}}^{\text{lower}}$ and $S_{\text{ID}}^{\text{upper}}$, this simplifies to: $w_{P}B<w_{\mathcal{N}}(C-c)$.
By selecting $w_{P}$ and $w_{\mathcal{N}}$ to satisfy this condition, we ensure that the scoring function $S_{\text{ID}}$ can differentiate between ID and OOD samples based on semantic similarity. ∎
Implications. This lemma establishes a theoretical basis for using the scoring function to distinguish ID from OOD samples by focusing on top-$k$ similarities, enhancing robustness by minimizing the effect of weakly relevant labels.

### v2

# CMOOD: Concept-based Multi-label OOD Detection

@@ Abstract
How can models effectively detect out-of-distribution (OOD) samples in complex, multi-label settings without extensive retraining? Existing OOD detection methods struggle to capture the intricate semantic relationships and label co-occurrences inherent in multi-label settings, often requiring large amounts of training data and failing to generalize to unseen label combinations. While large language models have revolutionized zero-shot OOD detection, they primarily focus on single-label scenarios, leaving a critical gap in handling real-world tasks where samples can be associated with multiple interdependent labels. To address these challenges, we introduce CMOOD, a novel zero-shot multi-label OOD detection framework. CMOOD leverages pre-trained vision-language models, enhancing them with a concept-based label expansion strategy and a new scoring function. By enriching the semantic space with both positive and negative concepts for each label, our approach models complex label dependencies, precisely differentiating OOD samples without the need for additional training. Extensive experiments demonstrate that our method significantly outperforms existing approaches, achieving approximately 95% average AUROC on both VOC and COCO datasets, while maintaining robust performance across varying numbers of labels and different types of OOD samples.
Codehttps://anonymous.4open.science/r/COOD-162F
Introduction
@@
As machine learning models become essential in a range of real-world applications, out-of-distribution (OOD) detection has gained increasing importance (Yang et al. 2024). OOD detection is particularly critical in fields such as autonomous driving (Elhafsi et al. 2023), medical diagnostics (Huang et al. 2024), and surveillance (Sultani et al. 2018), where detecting data that deviates from the training distribution is crucial to prevent safety risks or incorrect decisions (Hendrycks et al. 2019b). Thus, developing robust OOD detection techniques is key to ensuring model reliability in unpredictable environments.
@@
Large Models for OOD Detection (LM-OOD). The rise of large models, particularly Vision-Language Models (VLMs) and MultiModal Large Language Models (MLLMs), has redefined the landscape of OOD detection (Xu and Ding 2024). Traditional OOD detectors, such as Maximum Softmax Probability (MSP) and Mahalanobis distance-based techniques (Lee et al. 2018), typically require extensive task-specific training on in-distribution (ID) data. In contrast, LLM-based methods leverage pre-trained knowledge, allowing for zero-shot and few-shot OOD detection. Models like CLIP (Radford et al. 2021) exemplify this shift by achieving effective OOD detection with minimal task-specific training, relying on robust, pre-trained representations from multimodal datasets. Recent advancements, such as NegLabel (Jiang et al. 2024), refine zero-shot OOD detection by introducing negative mining strategies to identify semantically meaningful OOD categories. This highlights that carefully selected negative labels can significantly improve detection performance without retraining. LM-OOD methods offer distinct advantages over traditional approaches, including strong performance in limited data, high adaptability across various tasks, and improved computational efficiency through reduced dependence on task-specific data preparation and retraining (Ming et al. 2022; Miyai et al. 2024; Xu and Ding 2024).
@@
Figure 1: Motivation for CMOOD. Traditional methods struggle with complex multi-label cases. Our approach expands the label space with positive and negative concepts, enabling robust detection of complex OOD samples like “Okapi" and “Spork".
@@
Limitations in Current LM-OOD. Large models for OOD detection are primarily designed for single-label tasks, where each input corresponds to a single, definitive label (Fort et al. 2021; Ming et al. 2022). However, many real-world applications are inherently multi-label in nature: medical imaging often requires identifying multiple coexisting conditions (Kermany et al. 2018; Huang et al. 2024). Current state-of-the-art methods like NegLabel (Jiang et al. 2024) encounter limitations in these multi-label contexts, as seen in Fig. 1: First, the complexity of semantic similarity computations increases significantly with multiple concurrent labels compared to single labels. Second, these methods assume OOD samples are semantically distant from ID classes, which may not hold in multi-label settings where novel combinations of known concepts could form valid OOD cases (Wang et al. 2023; Nie et al. 2024). Additionally, these approaches struggle to effectively model complex label co-occurrence patterns and conditional dependencies, both crucial in multi-label contexts (Zhang and Zhou 2014; Yu et al. 2014). These limitations highlight an essential gap in adapting LM-OOD methods for multi-label tasks, where it is necessary to differentiate between known and unknown label combinations within complex label spaces for practical application (Miyai et al. 2024; Yang et al. 2024).
@@
Our Proposal: Extending LM-OOD to Multi-Label Settings. In this work, we introduce CMOOD (concept-based OOD detection), the first multi-label OOD detection framework that leverages VLMs, specifically CLIP (Wang et al. 2023), in a zero-shot setting. CMOOD addresses the challenges of multi-label OOD detection by incorporating a concept-based label expansion strategy. It enriches the base label set with two fine-grained concepts: positive concepts, which capture complex semantic details related to ID classes, and negative concepts, which, filtered by a similarity threshold, introduce semantically distant features to strengthen ID-OOD separation. CMOOD embeds these expanded concept labels alongside the original base labels into a novel scoring function that accounts for both scenarios: OOD multi-label inputs that share more similarity with ID classes (e.g., Giraffe and Okapi) and those with less similarity to any ID class. This design enables precise detection of subtle distinctions between ID and OOD samples without requiring additional training. Our technical coßntributions are summarized as follows:
@@
  • Superior Performance and Efficiency: Our method significantly outperforms existing approaches on standard benchmarks, with a high throughput of  800 images per second on CLIP-B/16.
Proposed CMOOD Method
Figure 2: Overview of CMOOD. The Concept Generation module uses LLMs to expand base labels into positive ($\mathcal{P}$) and negative ($\mathcal{N}$) concept sets, enhancing the ID-OOD boundary. Positive concepts capture fine-grained, ID-aligned features, while negative concepts provide contrasting OOD-aligned features. The Similarity and ID Score Computation module encodes an input image and computes similarity scores. An ID score based on top-$k$ similarities then classifies the image for ID/OOD.
Preliminaries on OOD Detection
In multi-label OOD detection, we determine whether an input image $I\in\mathcal{I}$ is ID or OOD. Let $\mathcal{B}=\{\ell_{1},\ell_{2},\dots,\ell_{|\mathcal{B}|}\}$ denote the set of known ID classes, where each image $I$ can have multiple labels from $\mathcal{B}$. We define an ID scoring function $S_{\text{ID}}:\mathcal{I}\rightarrow\mathbb{R}$ with the decision following criteria
@@
where $\gamma$ is a threshold determined through validation, and $\mathcal{D}_{\text{ID}}$, $\mathcal{D}_{\text{OOD}}$ are the ID and OOD distributions, respectively.
Revisit VLM-based OOD Detection
VLMs bring powerful multimodal capabilities for OOD detection, allowing more flexible, context-aware identification of OOD samples (Miyai et al. 2024; Xu and Ding 2024). Large-scale pre-trained VLMs, such as CLIP (Radford et al. 2021) and GPT-4V (Zhang et al. 2023), enable models to process images and text prompts together, enhancing adaptability and precision across diverse visual and textual domains (Fort et al. 2021).
VLM-based OOD detection generally follows two primary approaches (Xu and Ding 2024): Prompting-based Detection, which directly prompts VLMs to respond with OOD indicators (Cao et al. 2023), and Contrasting-based Detection, which uses multimodal VLMs pre-trained with contrastive objectives to distinguish OOD samples (Ming et al. 2022). This work focuses on the contrasting-based approach in multimodal contexts, as it is well-suited for enhancing OOD detection by amplifying distinctions between ID and OOD classes. Methods like NegLabel (Jiang et al. 2024) and NegPrompt (Li et al. 2024a) fall under this category. NegLabel enhances OOD detection by introducing negative labels to contrast with ID classes, while NegPrompt uses learned negative prompts to emphasize OOD differences by contrasting them with ID prompts.
@@
where $h$ is the embedding of $x$; $e_{i}$ and $e_{j}^{-}$ are embeddings for ID and negative representations (NegPrompt (Li et al. 2024a), NegLabel(Jiang et al. 2024)) from sets $Y$ and $Y^{-}$, respectively; $\text{sim}(h,e)$ denotes the similarity (e.g., cosine similarity) between embeddings. The numerator measures similarity to ID representations, while the denominator amplifies the detection signal with both ID and negative similarities. Note that our definition is a little different: a smaller $S(x)$ means a larger probability of $x$ being an OOD sample.
@@
Moreover, extending single-label OOD detection methods (Sun et al. 2021; Liu et al. 2020) to multi-label contexts presents additional challenges. Multi-label settings often involve significant semantic overlap between labels and co-occurring ID and OOD labels, complicating the extraction of effective negative labels or prompts. These issues highlight the need for advanced detection techniques to capture the complex dependencies and relationships inherent in multi-label OOD detection tasks.
Overview of CMOOD Method
In fact, it is not new to use LLM to assist in generating descriptions and then using CLIP models for classification tasks like DCLIP (Menon and Vondrick 2022). However, there are two problems with the application of this method: (1) the generated description only contains texts with similar semantic meanings, and the description of the text space is not comprehensive. (2) When considering the classification task, the OOD detection task does not bring improvement, and even brings negative benefits as show in Table 1.
The CMOOD framework improves OOD detection by refining the decision boundary between ID and OOD samples using fine-grained concepts. By introducing positive and negative concept sets, $\mathcal{P}$ and $\mathcal{N}$, CMOOD leverages LLMs to add contextual information around ID samples, thereby enhancing sensitivity to OOD cases.
CMOOD is built on two core components (see Fig. 2):
@@
    Concept Generation and Similarity Measure: To distinguish ID from OOD samples, we generate positive concepts ($\mathcal{P}$) and negative concepts ($\mathcal{N}$) that are closely and distantly related to ID classes, respectively. Positive concepts, mined using LLMs, capture domain-relevant features of ID samples at a fine-grained level, while negative concepts are chosen based on their semantic distance from ID classes to enhance contrast.
@@
    Similarity and ID Score Computation: For each input image $I$, we compute similarity scores with these concept sets and the base label set $\mathcal{B}$, forming a robust semantic space for evaluating ID-OOD relationships. After this, we define an ID score $S_{\text{ID}}(I)$ that aggregates the image’s alignment with base concepts ($\mathcal{B}$) and contrasts it with the positive and negative concepts ($\mathcal{P}$ and $\mathcal{N}$). When this score falls below a predefined threshold $\gamma$, the image is classified as OOD. This scoring method sharpens the decision boundary by leveraging contrasts with positive and negative concepts giving more accurate OOD detection.
The pseudocode in Algo. 1 outlines the full procedure for CMOOD. To formalize the OOD detection decision, we compute an ID score and classify an image $X$ as OOD if this score falls below a threshold $\gamma$. Specifically, the decision function $\tilde{Y}$ is defined as follows:
@@
Advantages. The proposed CMOOD approach is computationally efficient, leveraging pre-trained embeddings without requiring additional model training. By incorporating positive and negative concepts, the model gains enhanced semantic understanding, allowing for clearer differentiation between ID and OOD samples. The focus on top-$k$ similarity values makes the method robust to noise, ensuring stable performance across varied datasets and label sets.
Algorithm 1 Multi-Label OOD Detection with CMOOD
0: Image II, base label set $\mathcal{B}$, threshold $\gamma$, top-$k$ parameter $k$
0: Classification of II as ID or OOD
1:Concept GenerationConcept Generation): Generate positive concepts $\mathcal{P}$ and negative concepts $\mathcal{N}$ using an LLM.
@@
3:Top-$k$ Similarity CalculationSimilarity and ID Score Computation): Calculate the top-$k$ mean similarity scores, $\mu_{k}(\mathcal{B},I)$, $\mu_{k}(\mathcal{P},I)$, and $\mu_{k}(\mathcal{N},I)$.
4:ID Score ComputationSimilarity and ID Score Computation): Using the top-$k$ similarities, compute the ID score $S_{\text{ID}}(I)$ according to Eq. (7).
5:Decision: if $S_{\text{ID}}(I)>\gamma$ then classify II as ID; else classify II as OOD.
Concept Generation
Motivation. The concept generation process in CMOOD creates a rich semantic space that strengthens ID-OOD distinctions. By constructing a set of positive concepts, $\mathcal{P}$, aligned with ID samples, and a set of negative concepts, $\mathcal{N}$, that enhance the contrast with OOD samples, we can refine the decision boundary between ID and OOD. The sets $\mathcal{P}$ and $\mathcal{N}$ are complementary: $\mathcal{P}$ effectively captures multi-label inputs that share similarities with ID classes, while $\mathcal{N}$ captures multi-label inputs whose components are all dissimilar from ID samples. Together, these complementary concept sets enable CMOOD to model both subtle distributional shifts and drastic deviations, leading to improved OOD detection across varying degrees of distribution shift.
Positive Concept Mining $\mathcal{P}$. To build a comprehensive set of positive concepts that accurately represents ID characteristics, we use a two-stage approach inspired by LF-CBM (Oikarinen et al. 2023). This process enriches each base label by capturing specific attributes that reinforce its identity within the ID class.
@@
    Concept Querying: We prompt GPT-4 to generate concepts in three distinct contexts for each target object, using structured prompts tailored to elicit three categories: features (promptF), superclasses (promptS), and commonly associated items (promptC). This contextual querying ensures that each concept captures detailed, complex information, improving the clarity and consistency of the concept pool. Each prompt is crafted to encourage concise, relevant responses without qualifiers, which enhances precision, detailed in Supplementary Material §3.1.
@@
    Concept Filtering: After generating these candidate concepts, we apply filtering criteria to retain only the most distinctive and relevant features. This step is essential to ensure that the concept set effectively captures the ID characteristics required to differentiate ID from OOD. Our filtering process, detailed in the Supplementary Material §3.1, refines the generated concepts into a cohesive and meaningful set of positive labels.
@@
Negative Concept Mining $\mathcal{N}$. To reinforce the distinction between ID and OOD samples, we develop a set of negative concepts, $\mathcal{N}$, using a process called NegMining (Jiang et al. 2024). This approach helps us identify concepts that are semantically distant from ID labels, forming a contrasting boundary that enhances OOD detection.
@@
$\mathcal{N}$ $=\left\{n\mid\text{Sim}(\tilde{n}_{i},\mathcal{B})<\tau_{i}\right\},$ (6)
@@
By expanding the base label set $\mathcal{B}$ to include positive concepts $\mathcal{P}$ and contrasting them with negative concepts $\mathcal{N}$, we create an enriched semantic space. This structure allows for more precise discrimination between ID and OOD samples, capturing both the core characteristics of the ID classes and intra-class and inter-class similarity effectively.
Figure 3: t-SNE figure of label and concept text embeddings, along with corresponding image examples. In scenarios with multiple labels and objects, it is difficult to model the OOD detection problem using a single similarity measure. Instead, the COOD method is employed to address this issue.
Similarity and ID Score Computation
After generating our positive and negative concepts, CMOOD computes an ID score for each input image $I$, based on its similarity with the positive concept set $\mathcal{P}$, negative concept set $\mathcal{N}$, and base labels $\mathcal{B}$. This score helps determine if the image exhibits more ID or OOD characteristics. As shown in Figure 3, the positive concepts are distributed around the base labels, while the negative concepts are located at a relatively distant position in the text embedding space, showing a significant distinction. Therefore, to capture this distinction, we model the problem by adjusting the base label-image similarity through two components: (1) Subtracting the intra-class similarity between nearby positive concepts and the image, which captures within-distribution variations. Positive concepts are useful when identifying multi-label OOD inputs where it has some similarities with ID class like Giraffe and Okapi in Fig. 1. (2) Subtracting the inter-class similarity between distant negative concepts and the image, which captures cross-distribution relationships. This is to penalize the score when multi-label OOD input is not relevant to ID class.
ID Score Definition. The ID score $S_{\text{ID}}(I)$ for an image $I$ is:
$S_{\text{ID}}(I)=\underbracket{\left[\mu_{k}(\mathcal{B},I)-\mu_{k}(\mathcal{P},I)\right]}_{S_{\text{A}}}+\underbracket{\left[\mu_{k}(\mathcal{B},I)-\mu_{k}(\mathcal{N},I)\right]}_{S_{\text{B}}}$ (7)
where $\mathcal{B}$ are the base labels, $\mathcal{P}$ are the positive concepts, and $\mathcal{N}$ are the negative concepts.
In general, when an out-of-distribution (OOD) input has some similarity to an in-distribution (ID) class, we penalize $S_{A}$, whereas when an OOD input has no similarity to any ID class, we penalize $S_{B}$. However, for ID inputs, both $S_{A}$ and $S_{B}$ are larger. Consequently, ID inputs receive higher scores, while OOD inputs receive lower scores. Please refer to the theoretical bounds for details in Supplementary Material §2 and Fig. 1 that validate the scoring function’s effectiveness.
@@
Interpretation of the ID Score. The scoring function $S_{\text{ID}}(I)$ captures the degree of alignment of an image within ID or with OOD characteristics. For ID samples, we anticipate high alignment with base labels $\mathcal{B}$ , yielding a high $S_{\text{ID}}(I)$. Conversely, OOD samples will likely align more closely with negative concepts $\mathcal{N}$, resulting in a lower ID score. By focusing on the top-$k$ similarities, this approach enhances robustness to noise and emphasizes the most semantically relevant features.
This mechanism refines the boundary between ID and OOD samples by leveraging positive and negative contrasts to base labels, resulting in more accurate and reliable OOD detection.